An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed with a mean of 800 hours and a standard deviation 40 hours. Test the null hypothesis that =800 hours against the alternative ≠800 hours if a random sample of 20 bulbs have an average of 700 hours. Use alpha = 0.05 level of significance.

1
Expert's answer
2021-12-01T11:20:32-0500

The following null and alternative hypotheses need to be tested:

"H_0: \mu=800"

"H_1:\mu\not=800"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:



"z=\dfrac{x-\mu}{\sigma/\sqrt{n}}=\dfrac{788-800}{40/\sqrt{30}}\approx-1.643"

Since it is observed that"|z|=1.643<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-1.643)=0.100383," and since "p=0.100383>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 800, at the "\alpha=0.05" significance level.